OpenStudy (nuccioreggie):

Below are the data collected from two random samples of 500 American adults on the number of hours they work per day (rounded to the nearest hour): Number of hours of work per day 6 7 8 9 10 Sample A: Number of adults 60 90 145 150 55 Sample B: Number of adults 70 80 140 145 65 Ryan concludes that adults spend a mean of 8 hours working each day. Malia thinks the mean is 9 hours. Who is correct—Ryan or Malia? Explain your answer in two or three sentences. please EXPLAIN the ANSWER

OpenStudy (anonymous):

yes 8hrs so ryan is correct

OpenStudy (nuccioreggie):

please explain in to senteces

OpenStudy (michele_laino):

we have to compute the mean of both groups of data

OpenStudy (nuccioreggie):

ok

OpenStudy (anonymous):

There are 2 samples of 500 people each → that's 1000 people altogether who responded to the survey. Add up the total number for each response, so 60 + 70 = 130 people said they work 6 hours. Then 90 + 80 = 170 people said they worked 7 hours, and so on. The mean is each person's response divided by the total amount of responses. It would take too long to add up 6 hours 130 times, so we multiply 6*130 and do the same for each response. \[\frac{ 6(60+70)+7(90+80)+8(145+140)+9(150+145)+10(55+65) }{ 1000 }\]